OpenStudy (anonymous):

The taxi cost $2.25 for the first mile plus 15 Cents for each additional tenth of a mile. For a 5.2-mile trip, Eric paid $10 and told the driver to keep the change as a tip. How Much Was The Drivers Tip?? Please Help!!! Thanks SO Much!!

OpenStudy (agl202):

5.2 - 1 = 4.2 .15*42 = $6.30 + $2.25 = $7.55 10-7.55 = $2.45 for tips.

OpenStudy (agl202):

Medal?